# EDGE: Unknown-aware Multi-label Learning by Energy Distribution Gap Expansion

### v1

@@ Introduction
Traditional classification methods mainly focus on In-Distribution (ID) scenarios, where the training set and the test set share the same data distribution. However, in practical applications, it is unavoidable that some test samples are Out Of the Distribution (OOD) of the training ones (Yang et al. 2021). In this case, an ID model will classify the OOD samples into one of the ID classes due to the notorious overconfidence issue. Such wrong predictions seriously degenerate the model performance. In view of this, OOD detection has attracted rising attention in recent years, intending to discriminate the OOD samples from the ID ones (Huang and Li 2021; Wang et al. 2022c; Djurisic et al. 2023)
@@
Multi-label classification (MLC) focuses on the instances that are simultaneously assigned with multiple labels (Liu et al. 2022). This topic mainly falls into two categories: problem transformation and algorithm adaptation. For the former, most studies tend to simplify MLC into multi-class problems (Jernite, Choromanska, and Sontag 2017; Wallach et al. 2019). In contrast, algorithm adaptation leverages traditional machine-learning methods to model data (Gong et al. 2014; Wang et al. 2016; Chen et al. 2019). Different from traditional MLC work, this paper dedicates to exploring OOD detection in multi-label learning, which is still an underexplored problem.
@@
Traditional long-tailed learning advocates re-balancing imbalanced distributions with undersampling (Peng et al. 2019), oversampling (Mullick, Datta, and Das 2019), or re-weighting (Menon et al. 2021; Alshammari et al. 2022). However, these methods proved to be inapplicable in (1) OOD detection (Wang et al. 2022b) and (2) multi-label settings (Duarte, Rawat, and Shah 2021), which could not be simply transferred to long-tailed OOD detection. Also, the research for imbalanced multi-label learning (Guo and Wang 2021; Wu et al. 2020) is not well-suitable. Fortunately, Wang et al. (Wang et al. 2022b) propose the first learning framework to instruct OOD detection on multiclass long-tailed scenarios. Inspired by this work, this paper attempts to explore the study of long-tailed OOD detection for multi-label scenarios.
@@
Enlarge Distribution Distance. The reduction in logits for OE samples allows for the accurate identification of most OOD samples. Yet, a subset of challenging cases persists among tail ID samples and semantically similar near-boundary OOD samples (Wang et al. 2023). Our analysis suggests that this problem stems from a critically narrow gap in energy distribution, as depicted in Figure 1 (a). The considerable overlap in distribution areas culminates in ambiguous distinctions. A type of partial contrastive learning (Wang et al. 2022b) proves to be a potent method for distancing tail ID samples from their OOD counterparts. Nonetheless, its assumption of class-conditional independence is untenable for the unknown joint distributions encountered in multi-label settings. In response to these problems, we opt to treat the $k$ ID samples with the lowest scores collectively in contrast with all OE sample, and propose an energy-based loss function $\mathcal{L}_{\textup{gap}}$ to expand the energy gap:
@@
It has been proved that a certain amount of auxiliary out-of-distribution data would extremely boost the discrimination of OOD detectors (Hendrycks, Mazeika, and Dietterich 2019). In this case, the model establishes a sufficiently strong decision boundary by previewing information from the real distribution. However, the detection effect is uncertain when using different $\mathcal{D}_{\textup{OE}}$ for training, which indicates that the success of outlier exposure is quite dependent on proper auxiliary data. Thus, a key challenge in our framework is how to select a more reliable auxiliary dataset? By observing the feature space, it could be found the samples with similar semantic information usually assemble (Du et al. 2022). This trivial nature results in a serious failure to distinguish the samples with strong semantic correlation. Thus, we argue that these OOD samples could play a decisive role in improving the discrimination of the model. Based on this discovery, we propose a reasonable feature-based dilation distance to measure the suitability of outlier samples:
@@
Overall Performance. We present the overall comparative results on various ID and OOD datasets in Table 3. We get some observations below: (1) Our proposed EDGE significantly outperforms all comparative methods; (2) Our EDGE achieves a significant improvement in OOD detection with a less loss of ID performance, promising a relatively balanced result between known and unknown domains; (3) The complexity of the $\mathcal{D}_{\textup{in}}$ and $\mathcal{D}_{\textup{OE}}$ directly affects both ID classification and OOD detection. In larger-scale NUS-WIDE, most comparative methods produce serious degeneration, especially on iSUN, where this dataset contains some similar scene images, like the jungle, with the images in NUS-WIDE. However, our method still maintains promising performance while resisting hard samples; (4) An interesting point is that both OE and EnergyOE achieve an unstable performance, even poorer than some post-hoc methods. It indicates that simply transferring general outlier exposure methods is not always feasible; (5) OE and EnergyOE perform better when utilizing MaxEnergy while not JointEnergy. A reasonable explanation is that traditional methods are applied in multi-class scenarios and prefer to identify OOD samples with their maximum logit. This nature makes them function poorly in multi-label scenarios.

### v2

@@ Introduction
Traditional classification methods mainly focus on In-Distribution (ID) scenarios, where the training set and the test set share the same data distribution. However, in practical applications, it is unavoidable that some test samples are Out Of the Distribution (OOD) of the training ones (Yang et al. 2021). In this case, an ID model will classify the OOD samples into one of the ID classes due to the notorious overconfidence issue. Such wrong predictions seriously degenerate the model performance. Given this, OOD detection has attracted rising attention in recent years, intending to discriminate the OOD samples from the ID ones (Huang and Li 2021; Wang et al. 2022c; Djurisic et al. 2023)
@@
Multi-label classification (MLC) focuses on the instances simultaneously assigned with multiple labels (Wang et al. 2024; Liu et al. 2022; Huang et al. 2020). This topic mainly falls into two categories: problem transformation and algorithm adaptation. For the former, most studies tend to simplify MLC into multi-class problems (Jernite, Choromanska, and Sontag 2017; Wallach et al. 2019). In contrast, algorithm adaptation leverages traditional machine-learning methods to model data (Gong et al. 2014; Wang et al. 2016; Chen et al. 2019). Unlike traditional MLC work, this paper is dedicated to exploring OOD detection in multi-label learning, which is still an underexplored problem.
@@
Traditional long-tailed learning advocates re-balancing imbalanced distributions with undersampling (Peng et al. 2019), oversampling (Mullick, Datta, and Das 2019), or re-weighting (Alshammari et al. 2022; Wang et al. 2023b; Yang et al. 2024). However, these methods proved to be inapplicable in (1) OOD detection (Wang et al. 2022b) and (2) multi-label settings (Duarte, Rawat, and Shah 2021), which could not be simply transferred to long-tailed OOD detection. Also, the research for imbalanced multi-label learning (Guo and Wang 2021; Wu et al. 2020) is not well-suitable. Fortunately, Wang et al. (Wang et al. 2022b) propose the first learning framework to instruct OOD detection on multiclass long-tailed scenarios. Inspired by this work, this paper attempts to explore the study of long-tailed OOD detection for multi-label scenarios.
@@
Enlarge Distribution Distance. The reduction in logits for OE samples allows for the accurate identification of most OOD samples. Yet, a subset of challenging cases persists among tail ID samples and semantically similar near-boundary OOD samples (Wang et al. 2023a). Our analysis suggests that this problem stems from a critically narrow gap in energy distribution, as depicted in Figure 1 (a). The considerable overlap in distribution areas culminates in ambiguous distinctions. A type of partial contrastive learning (Wang et al. 2022b) proves to be a potent method for distancing tail ID samples from their OOD counterparts. Nonetheless, its assumption of class-conditional independence is untenable for the unknown joint distributions encountered in multi-label settings. In response to these problems, we opt to treat the $k$ ID samples with the lowest scores collectively in contrast with all OE sample, and propose an energy-based loss function $\mathcal{L}_{\textup{gap}}$ to expand the energy gap:
@@
It has been proved that a certain amount of auxiliary out-of-distribution data would extremely boost the discrimination of OOD detectors (Hendrycks, Mazeika, and Dietterich 2019). In this case, the model establishes a sufficiently strong decision boundary by previewing information from the real distribution. However, the detection effect is uncertain when using different $\mathcal{D}_{\textup{OE}}$ for training, which indicates that the success of outlier exposure is quite dependent on proper auxiliary data. Thus, a key challenge in our framework is how to select a more reliable auxiliary dataset? By observing the feature space, it could be found the samples with similar semantic information usually assemble (Du et al. 2022). This trivial nature results in a serious failure to distinguish the samples with strong semantic correlation. Thus, we argue that these OOD samples could play a decisive role in improving discrimination. Based on this discovery, we propose a reasonable feature-based dilation distance to measure the suitability of outlier samples:
@@
Overall Performance. We present the overall comparative results on various ID and OOD datasets in Table 3. We test various inference methods as well as different outlier exposure methods under the same training criterion and pre-trained model. We get some observations below: (1) Our proposed EDGE significantly outperforms all comparative methods; (2) Our EDGE achieves a significant improvement in OOD detection with a less loss of ID performance, promising a relatively balanced result between known and unknown domains; (3) The complexity of the $\mathcal{D}_{\textup{in}}$ and $\mathcal{D}_{\textup{OE}}$ directly affects both ID classification and OOD detection. In larger-scale NUS-WIDE, most comparative methods produce serious degeneration, especially on iSUN, where this dataset contains some similar scene images, like the jungle, with the images in NUS-WIDE. However, our method still maintains promising performance while resisting hard samples; (4) An interesting point is that both OE and EnergyOE achieve an unstable performance, even poorer than some post-hoc methods. It indicates that simply transferring general outlier exposure methods is not always feasible; (5) OE and EnergyOE perform better when utilizing MaxEnergy while not JointEnergy. A reasonable explanation is that traditional methods are applied in multi-class scenarios and prefer to identify OOD samples with their maximum logit. This nature makes them function poorly in multi-label scenarios.
@@ Conclusion
Acknowledgments
This work was supported in part by the National Key R&D Program of China under Grant 2018AAA0102000, in part by National Natural Science Foundation of China: 62236008, U21B2038, U23B2051, 61931008, 62122075, 62206264, and 92370102, in part by Youth Innovation Promotion Association CAS, in part by the Strategic Priority Research Program of the Chinese Academy of Sciences, Grant No. XDB0680000, in part by the Innovation Funding of ICT, CAS under Grant No.E000000, in part by the China National Postdoctoral Program for Innovative Talents under Grant BX20240384.